Cost of Preferred Stock with Flotation Costs

Burnwood Tech plans to issue some $60 par preferred stock with a 6% dividend. A similar stock is selling on the market for $61. Burnwood must pay flotation costs of 5% of the issue price. What is the cost of the preferred stock? Round your answer to two decimal places.
Answer #1

Solution :

Fixed dividend = 60 × 6% = $ 3.6

Net proceeds = market price - flotation cost = 61 - (5% of 61) = $57.95

Cost of preferred stock = fixed dividend / net proceeds = 3.6/57.95= $6.21%